40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Mar 11 18:43:07 2018, Craig, DailyProgress, PEM, Notes from today's work 
    Reply  Mon Mar 12 01:25:23 2018, Craig, DailyProgress, PEM, Notes from today's work 
       Reply  Mon Mar 12 15:22:42 2018, Craig, Jaime, DailyProgress, PEM, Notes from today's work 
    Reply  Thu Mar 15 10:29:04 2018, awade, DailyProgress, scatter, Addressing 500 Hz scatter pickup 2018-03-14_16.58.21.jpg2018-03-14_20.40.18.jpg2018-03-14_16.58.13.jpg
Message ID: 2135     Entry time: Thu Mar 15 10:29:04 2018     In reply to: 2129
Author: awade 
Type: DailyProgress 
Category: scatter 
Subject: Addressing 500 Hz scatter pickup 

Going back the original issue of scattering, it appears that there is light being back reflected from somewhere in the post PMC path but before the reference cavities.  

Reducing number of optics after the North PMC 

I had installed a bunch of polarization optics before the north 14.75 MHz EOM in an effort to reduce RFAM (see attachement 1).  It looks like stuffing so many optics in such a small space is a bad idea.  You can see weak retro reflected beams from the wave plates and, probably, the PBS as well.  The short propagation distance makes it difficult to angle optics enough to be able to separate them from the main beam laterally to dump.  The EOM can't really be moved because the mode matching solution is a little tight for the available space.

After talking with rana and Craig yesterday it seems like the Pre Mode Cleaner (PMC) should be filtering polarization well enough when locked that the PBS and quarter-wave plate (QWP) are unnecessary. I removed all but the half-wave plate (HWP) and checked the residual polarization on transmission with a diagnostic PBS in place. I found 2 µW of power out of 1.2 mW was remaining when tuned all the way to s-pol.: this is a 1:600 extinction ratio which is about what we would expect from such a beam cube.  This measurement may be biased by the lower limit of the power meter, PBS should be giving 1:1000.  

I moved the PBS to before the PMC to clean up light out of the 21.5 MHz PMC phase modulator. The only optics in the post PMC-> EOM path are now a lens, a steering mirror and a half-wave plate (see attachment #2).  After realigning the PMC cavity and the north refcav I was able to reduce the RFAM to -55 dBm, which is good enough for now.  These slight changes in RFAM level mean that the FSS offset will need some adjustment.  I was unable to see any improvement in the beat spectrum as the beat note had drifted down to 2 MHz.  I turned the heating down a small amount and left it overnight to settle.

I didn't angle the HWP or lens by that much, this shouldn't be necessary because the PMC is a traveling wave cavity.  The elements should be pretty close to normal.  The glass beam dump should be checked to ensure it is not clipping any retro-reflected beams on the rough edge of the glass.

Clamping down the PMC

I never clamped down the PMC. It is just sitting on the ball baring points. This isn't great.

When I realized the tapped holes on the side of the base I went looking for clamps.  They are pictured in attachment #3 but they do not fit.  It turns out there were some issues with the choice of ball bearings on which the PMC sits.  The ball barrings sit over holes so that the PMC when placed will realigned exactly with its previous position on the base.  Antonio had found that the holes drilled for the ball barrings were spec'ed a little too big.  For standard increments of bearings size the closest size fits nicely over the hole but under force they actually slip down into the hole and are almost impossible to get out.  He bought the next ball bearing size up. However, this means that the clamps no longer reach the full height PMC assembly.  The assumed tolerances were made too tight on all these components, the next edition of drawings should allow for some wiggle room.

The drawings should be updated with at least 1-3 mm of range on slot cut side pieces for the clamps so there is room for changes in height due to ball bearing size.  Possibly even more, if future people want to put Viton or Sorbothane dampening into the clamping. The non-tapped holes should also be changed to through-all. Or at least drill with a narrower diameter through-all. This will help future users poke out objects that get stuck in the holes.  

For that matter the design of the clamps seems wrong.  There is a bar that goes over the top that is fixed with a slot-cut piece affixed to each side. This is intuitively wrong as the bolts all go in horizontally when the clamping force needs to be applied is downwards!  It means that the clamps are locking a vertically applied force from the sides; to bolt the PMC down you need to apply force to the bar and tighten the bolts at the same time for two different clamping bars.  The screws should have at least one vertical pair on each clamp so that tension can be applied in the same direction as the clamping force.  

PMC documents on the DCC

For future reference, here is a list of all the PMC documents on the DCC:

Evan's technical note for PMC design considerations: LIGO-T1600071.

I can't find assembly procedures on the DCC.  There was a report from one of Kate Dooley's summer students, LIGO-T1600503, that shows a jig for gluing the PZT. 

 

Quote:

Today I buzzed the table and determined there was a strong 500 Hz dirty resonance on the first steering mirror after the PMC. 
This caused me to go around tightening bolts everywhere, including the offending steering mirror and the optics around it.  This did not reduce the resonance.
I tightening the PMC REFL steering mirror as well, and this caused a misalignment onto the PMC REFL PD.  This took me a little while to figure out why the North path refused to lock.  I realigned the PMC REFL steering mirror into the PD.
After I got the North PMC locking again, the North path itself was not locking anymore.  I reranged the autolocker slow volts, but this did not help. 
Turns out the North Trans PD threshold voltage was not high enough.  This is likely because of the bolt tightening, causing some slight misalignment into the North cavity, lowering the overall circulating power in the cavity.  I lowered the autolocker threshold from 1.1 volts to 1.0 volts, and aligned the North Trans PD.  We need to rescan the North cavity to get better alignment/mode matching, but I'm gonna put this off until we replace this offending 500 Hz post-PMC steering mirror.
While I was realigning the Trans PD, I noticed that even touching the trans optics tables causes large ~1Hz oscillations in the trans voltage.  This is definitely exacerbating any scattering problem we have.  Also, the Trans PD output for both paths is "breathing", going up and down with a period of about a minute.  This is bad for our autolocker's threshold.  It's possible that we should build two periscopes for the north and south paths to eliminate these elevated tables which cause coherent oscillations on all trans optics.  We could copy Tara's front periscope design.

 

Attachment 1: 2018-03-14_16.58.21.jpg  3.178 MB  Uploaded Thu Mar 15 14:11:10 2018  | Hide | Hide all
2018-03-14_16.58.21.jpg
Attachment 2: 2018-03-14_20.40.18.jpg  3.234 MB  Uploaded Thu Mar 15 14:11:27 2018  | Hide | Hide all
2018-03-14_20.40.18.jpg
Attachment 3: 2018-03-14_16.58.13.jpg  1.493 MB  Uploaded Thu Mar 15 14:11:43 2018  | Hide | Hide all
2018-03-14_16.58.13.jpg
ELOG V3.1.3-